### v1

@@ 5 Observability with SKA
Firstly, we compare our predictions with current upper limits from a number of facilities including MWA, LOFAR, GMRT, PAPER and HERA (we focus at where most of the measurements have been taken).
In Fig. 13 we show the 21cm power spectrum at = 10, 9, 8 and 7 for all the four models together with the available upper limits (see label). All our models are below these upper limits. However, the moderate and extreme Pop. III models are closer to the current HERA constraints at = 8 suggesting that soon these models can be either detected or ruled out.
@@
The sensitivity greatly improves for an observational campaign of 1000 hours (Fig. 16). In this case, models accounting for Pop. III early heating which is stronger than Pop. II are clearly distinguishable from models with no or mild Pop. III contribution. At = 10, 9 and 8, it is even possible to distinguish between intermediate and extreme Pop. III models at all scales within and , indicating that SKA observations of the 21cm power spectrum during the EoR are sensitive to the IGM heating occurred at . At = 7, the difference between the various models are less evident but intermediate and extreme Pop. III models still fall outside the uncertainty regions of weak Pop. III and No Pop. III. This indicates that models with a harder Pop. III X-ray emission can be disentangled from models with no or mild Pop. III X-ray emission even at = 7. Finally, we highlight that the impact of Pop. III stars is visible on the power spectrum as long as their X-ray contribution is boosted compared to Pop. II. The differences between the weak Pop. III and the No Pop. III scenario are so small they will not be visible even with 1000 SKA hours. This demonstrates that an upcoming 21cm power spectrum detection will shed light on the properties of the first sources establishing their contribution to the heating of the IGM.
@@ 6 Conclusions
In this work we investigated how the early () heating of the IGM due to Pop. III stars impacts the 21cm power spectrum during the EoR (z ) which will be observable by the SKA. We developed a framework of scaling relations between Pop. III star formation rate in mini-halos and the density field within the semi-analytical model meraxes that allows us to account for the radiative contribution of mini-halos that cannot be directly resolved in a large (L = 310 cMpc) cosmological simulation. The scaling relations are calibrated on a smaller and higher resolution simulation able to resolve all the mini-halos. We then investigated three Pop. III models each with different IMF, star formation efficiency and specific X-ray luminosity. This formalism is both based on realistic galaxy population and it provides an accurate estimation of the Pop. III backgrounds yielding a computation of the ionization and thermal state of the IGM which is crucial to obtain a reliable estimation of the 21cm signal and power spectrum during the EoR ( ). The key results can be summarized as follows:
@@ Acknowledgements
This research was supported by the Australian Research Council Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D, project #CE170100013). This work was performed on the OzSTAR national facility at Swinburne University of Technology. OzSTAR is funded by Swinburne University of Technology and the National Collaborative Research Infrastructure Strategy (NCRIS). This research was also undertaken with the assistance of resources from the National Computational Infrastructure (NCI Australia), an NCRIS enabled capability supported by the Australian Government. SB acknowledges support from grant PID2022-138855NB-C32 funded by MICIU/AEI/10.13039/501100011033 and ERDF/EU, and project PPIT2024-31833, cofunded by EU–Ministerio de Hacienda y Función Pública–Fondos Europeos–Junta de Andalucía–Consejería de Universidad, Investigación e Innovación. YQ is supported by the ARC Discovery Early Career Researcher Award (DECRA) through fellowship #DE240101129. Finally, EMV thank A. Mesinger and J. B. Munoz for insightful discussions regarding the development of the scaling relation model and C. Power for assistance with the N-body simulation.

### v2

@@ 5 Observability with SKA
Firstly, we compare our predictions with current upper limits from a number of facilities including MWA, LOFAR, GMRT, PAPER and HERA (we focus at where most of the measurements have been taken). In Fig. 13 we show the 21cm power spectrum at = 10, 9, 8 and 7 for all the four models together with the available upper limits (see label). All our models are below these upper limits. However, the moderate and extreme Pop. III models are closer to the current HERA constraints at = 8 suggesting that soon these models can be either detected or ruled out. In our model the main effect of different Pop. III models is to change the timing and amplitude of the peaks in the 21cm power spectrum rather than introducing a specific feature. We note that we did not account for spatial variations in the velocity acoustic oscillations (see Section 2.1) which would introduce wiggles in the at scales Mpc-1 (Cruz et al. 2024a). While this effect can be important when focusing on the 21cm power spectrum during the coupling and heating epochs, these fluctuations are quickly washed out at (see Fig. 13 and Section 6C in Cruz et al. 2024a).
@@
The sensitivity greatly improves for an observational campaign of 1000 hours (Fig. 16). In this case, models accounting for Pop. III early heating which is stronger than Pop. II are clearly distinguishable from models with no or mild Pop. III contribution. At = 10, 9 and 8, it is even possible to distinguish between intermediate and extreme Pop. III models at all scales within and , indicating that SKA observations of the 21cm power spectrum during the EoR are sensitive to the IGM heating occurred at . At = 7, the difference between the various models are less evident but intermediate and extreme Pop. III models still fall outside the uncertainty regions of weak Pop. III and No Pop. III. This indicates that models with a harder Pop. III X-ray emission can be disentangled from models with no or mild Pop. III X-ray emission even at = 7. Also Gessey-Jones et al. (2025) focused on the possibility of detecting difference Pop. III models using 21cm observations. In their analysis they included both REACH (sensitive to the 21cm global signal) and SKA-Low confirming that an observation of hours with SKA-Low can constrain the Pop. III IMF due to their difference in the X-ray heating. Finally, we highlight that the impact of Pop. III stars is visible on the power spectrum as long as their X-ray contribution is boosted compared to Pop. II. The differences between the weak Pop. III and the No Pop. III scenario are so small they will not be visible even with 1000 SKA hours. This demonstrates that an upcoming 21cm power spectrum detection will shed light on the properties of the first sources establishing their contribution to the heating of the IGM.
@@ 6 Conclusions
In this work we investigated how the early () heating of the IGM due to Pop. III stars impacts the 21cm power spectrum during the EoR () which will be observable by the SKA. We developed a framework of scaling relations between Pop. III star formation rate in mini-halos and the density field within the semi-analytical model meraxes that allows us to account for the radiative contribution of mini-halos that cannot be directly resolved in a large (L = 210 h-1 cMpc) cosmological simulation. The scaling relations are calibrated on a smaller and higher resolution simulation able to resolve all the mini-halos. We then investigated three Pop. III models each with different IMF, star formation efficiency and specific X-ray luminosity. This formalism is both based on realistic galaxy population and it provides an accurate estimation of the Pop. III backgrounds yielding a computation of the ionization and thermal state of the IGM which is crucial to obtain a reliable estimation of the 21cm signal and power spectrum during the EoR ( ). The key results can be summarized as follows:
@@ Acknowledgements
We thank the anonymous referee for very constructive comments that help improve the quality of this paper. This research was supported by the Australian Research Council Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D, project #CE170100013). This work was performed on the OzSTAR national facility at Swinburne University of Technology. OzSTAR is funded by Swinburne University of Technology and the National Collaborative Research Infrastructure Strategy (NCRIS). This research was also undertaken with the assistance of resources from the National Computational Infrastructure (NCI Australia), an NCRIS enabled capability supported by the Australian Government. SB acknowledges support from grant PID2022-138855NB-C32 funded by MICIU/AEI/10.13039/501100011033 and ERDF/EU, and project PPIT2024-31833, cofunded by EU–Ministerio de Hacienda y Función Pública–Fondos Europeos–Junta de Andalucía–Consejería de Universidad, Investigación e Innovación. YQ is supported by the ARC Discovery Early Career Researcher Award (DECRA) through fellowship #DE240101129. Finally, EMV thank A. Mesinger and J. B. Munoz for insightful discussions regarding the development of the scaling relation model and C. Power for assistance with the N-body simulation.
@@ Appendix A Including the H2 self-shielding in meraxes
Appendix B Impact of Pop. III star formation efficiency and IMF
While the 21cm power spectrum at is mostly sensitive to the amount of X-ray photons from Pop. III remnants, changing the star formation efficiency and the shape of the IMF impacts the 21cm signal evolution during the coupling and heating epochs. In Fig. 18 and 19 we show the evolution of the 21cm global signal and power spectrum for the No Pop. III model (black line), Weak (grey line), Moderate (cyan line), high SFE (dashed grey line) and LogE (dashed red line) models. Compared to the Weak Pop. III model, increasing the Pop. III star formation efficiency, increases both the amount of Lyman- and X-ray photons (they are both linked to SFR), hence the absorption trough shifts towards higher- (). Considering instead a top-heavy IMF, increases just the amount of Lyman- photons hence the absorption trough becomes significantly deeper while still being at the same redshift compared to the weak Pop. III model. We note that this result is a consequence of the fact that we are considering the LX/SFR and the shape of the IMF as two independent parameters. If, as shown in Sartorio et al. (2023), the X-ray emissivity is linked to the shape of the IMF, the LogE model is not realistic as we would need to enhance the X-ray emissivity as well.
We highlight that both the star formation efficiency and the shape of the IMF alone do not significantly affect neither the 21cm global signal nor the power spectrum at so a future 21cm detection with SKA-low likely will not help to constrain these parameters if they are not linked to a significantly stronger X-ray emissivity.